OpenStudy (anonymous):

how to graph {x}=2y? {x} is the fractional part of x..i.e x=[x]+{x}

OpenStudy (watchmath):

start by doing the graph of {x}=2y on the interval [0,1) first

OpenStudy (anonymous):

in the interval [0,1) we have x=2y

OpenStudy (anonymous):

then what, watchmath?

OpenStudy (watchmath):

Then what happen in [1,2) ? :D

OpenStudy (anonymous):

i saw the graph in wolframalpha..but doesnt make sense!

OpenStudy (watchmath):

The value of {x} on [2,3) is also [0,1)

OpenStudy (anonymous):

in [1,2), we have x-[x]=1-1=0 for x=1, x-[x]=1.1-1=0.1..and so we have the smae graph x=2y

OpenStudy (anonymous):

so we have x=2y in every such interval:) am i right?

OpenStudy (watchmath):

Yes, to be precise line segment of the same slope

OpenStudy (anonymous):

thnx..:)